# CycleManip: Enabling Cyclic Task Manipulation via Effective Historical Perception and Understanding

### v1

@@ 1 Introduction
The ability of a robot to autonomously take care of various tasks in our daily life is a long-term goal of the computer vision and robotics community [black2024pi0, zhao2023act, james2020rlbench, liu2023libero]. One key observation is that many tasks in the household involve repetitive and cyclic actions, such as dispensing several pumps of syrup or shaking a bottle until the contents are mixed [al2022impact, holland2020automation, rane2017assembly]. These tasks require the robot to perform cyclic actions and stop at an expected time, which presents significant challenges: the robot needs to execute repetitive actions and complet them within the desired time frame according to the user command or task execution status. Previous work on robot manipulation has focused on general tasks [chi2025dp, liu2024rdt]; however, relatively little work has explored how to enable robots to perform cyclic tasks. Moreover, there is a lack of benchmarks that provide sufficient data and evaluation tools to assess the applicability of existing methods to cyclic tasks.
In this work, we explore this crucial task, CycleManip, the problem of robotic cyclic manipulation where a robot must accurately execute cyclic motions and stop at the correct moment. Such tasks are common in daily scenarios, but their inherently non-Markovian nature makes them challenging: the correct decision at any moment depends not only on the current observation, but also on the accumulated progress within the cycle. As a result, autonomous policies must model historical temporal dependencies and reason about the progression across multiple cycles[whitehead1995reinforcement].
However, traditional imitation policies typically rely on short observation windows for action prediction [black2024pi0, ze20243dp3, brohan2022rt, kim2024openvla, team2024octo, wang2024rise, rdt2], which leads to failure in cyclic manipulation tasks. This is because short observations across cycles often appear similar, causing the model to confuse its decisions as shown in Figure 2 (a). For example, in a task that requires the robot to shake a bottle five times, the visual observation after each shaking step remains nearly identical, making it difficult for such models to infer whether the robot should continue or stop, since they cannot track how many shaking cycles have already been completed. An intuitive remedy is to expand the observation horizon, but doing so is costly: encoding and fusing high-dimensional visual observations at every timestep increases computation and latency.
@@
To support our framework, we present a cycle-based manipulation benchmark that provides a diverse set of simulated tasks with automated data generation and model evaluation tools. The benchmark is built on the RoboTwin 2.0 simulation platform [chen2025robotwin2], where we incorporate configurable cyclic action parameters into the data collection pipeline, enabling the generation of demonstrations with arbitrary numbers of repetitions and corresponding instructions. In addition, we develop an automated evaluation system in which an attempt is considered successful only if the policy not only completes the manipulation task but also performs the correct number of cycles.
@@ 2.1 Robotic Manipulation
Robotic manipulation is a fundamental challenge in robotics and is considered an essential cornerstone for achieving Artificial General Intelligence (AGI). Current dominant methodologies, such as Imitation Learning (IL) [chi2025dp, ze20243dp3, zhao2023act] and Vision-Language-Action (VLA) models [zitkovich2023rt2, kim2024openvla, black2024pi0, intelligence2025pi05], excel at modeling complex data distributions. They effectively predict subsequent actions based on current observations, demonstrating strong performance in sequential tasks. However, these approaches falter in cyclic tasks. When faced with similar inputs, the model struggles to distinguish the current phase of the cycle, which can lead to it becoming trapped in infinite loops or terminating prematurely. To solve this problem, we propose the CycleManip framework to enable cyclic task manipulation in an end-to-end imitation manner by the effective perception and understanding of historical information.
@@ 2.2 Historical Modeling
Historical modeling is important for Large-Language-Models [wang2025MemoryLLM], Vision-Language-Navigation [zhang2025MemoryVLN], video generation [xiao2025worldmem] and robot manipulation [shi2025memoryvla]. The historical information is used to address the limitations of short-term observations, by retrieval [yu2025context, xiao2025worldmem], large kernels [gu2025long], memory caching [zeng2024poliformer, liu2025ttf]. Recent robot manipulation methods integrate historical visual information for visual memory [liu2025ttf, shi2025memoryvla, sridhar2025memer], with a primary focus on long-horizon tasks. In contrast, this paper addresses the more challenging domain of cyclic tasks and introduces a cost-aware sampling strategy to integrate historical low-overhead proprioceptive information.
@@ 2.3 Cyclic Task
Cyclic tasks are important in daily life and represent a critical challenge that needs to be addressed for robot deployment. The execution of cyclic tasks can significantly enhance efficiency [hausser2014experimental, hall1988cyclic] and be crucial in environments such as factories, laboratories [al2022impact, holland2020automation, rane2017assembly] and human-robot collaboration scenarios [kheirabadi2023human, keshvarparast2024collaborative]. Some traditional works implement cyclic tasks by control strategies, which may be limited by poor adaptability to dynamic environments. Recently, some deep learning-based methods have been attempted for simple cyclic tasks [yang2022learning]; however, these methods are either limited to simple tasks with fixed cyclic times [torne2025learning], limited to specific task scenarios [liang2020teaching, huang2019accurate], or rely on external auxiliary models [chen2025history]. In this paper, we propose a method that enables diverse cyclic manipulation tasks with cyclic time control capability, in an end-to-end manner without leveraging any additional auxiliary models.
@@ 3.2 Effective Historical Perception
For high-overhead observations (e.g., point clouds for 3D-based methods and RGB images for 2D-based methods), we adopt a heuristic sampling strategy that selects past frames with a longer observation horizon while keeping the overall number of frames $K_{\text{high}}$ consistent with the original version. Specifically, given the first frame as 0 and the current frame as $t$, we perform right-sided binary sampling to collect $0.5\cdot K_{\text{high}}$ frames. And we apply exponential sampling from the $K_{\text{high}}$ frame, collecting another $0.5\cdot K_{\text{high}}$ frames according to the rule $K_{\text{high}}-2^{k}$, where $k$ denotes the sampling index.
@@ 3.3 Effective Historical Understanding
Challenge of Historical Understanding. While expanding the observation horizon provides the model with broader perception, an overload of information introduces challenges in feature encoding and understanding. Relying only on imitation learning supervision typically fails to enable the model to truly understand the inherent temporal process of cyclic tasks, as shown in Figure 2 (b). For example, in a hammering task, the historical observations before each strike differ due to different motion histories, yet the supervision signal remains the same (conduct a hammering). This discrepancy may force the model to learn features that converge to a local optimum, making it difficult for the model to capture the temporal information necessary for cyclic behavior.
@@ 3.4 Framework Architecture
Given the user instruction lanlan and historical observation $\{o_{i}\}_{i=1}^{t}=\{o_{high},o_{low}\}_{i=1}^{t}$, we first employ a cost-aware sampling strategy to obtain sampled observation $\mathcal{H}_{h}\left(\{{o^{h}_{i}}\}_{i=1}^{t}\right)$ and $\mathcal{H}_{l}\left(\{{o^{l}_{i}}\}_{i=1}^{t}\right)$. The language feature $f_{lan}$ is encoded by CLIP encoder [radford2021clip], the high observation feature $f_{h}$ is encoded by a point encoder [ze20243dp3]. For low-overhead observations, we employ a Transformer encoder [vaswani2017attention] to obtain $f_{l}$, where the global features are derived from the CLS token, and local features are extracted using an MLP from the recent frames. Then we employ MLPs to fuse $f_{l}$ and $f_{h}$ obtain $f_{lh}$, which will be used for diffusion decision and auxiliary task prediction. Finally, we concatenate the language feature $f_{lh}$ and observation feature $f_{lh}$ as the condition feature for diffusion model. We employ film conditioning [peebles2023dit] to output the action prediction. The final loss function of the model is:
@@ 4 CycleManip Benchmark
To support our framework, we build a benchmark for cyclic manipulation tasks based on the RoboTwin 2.0 platform [chen2025robotwin2]. We collect 8 cyclic manipulation task environments for convenient data collection and policy evaluation, as shown in Figure 4. More details can be found in the supplementary material.
@@
block
hammering
bottle
shaking
roller
rolling
carrot
cutting
dual-knife
chopping
egg
beating
chemical
mixing
morse
tapping
@@
Table 1: Performance comparison on various cyclic manipulation tasks. Our method outperforms all baselines in Success Rate (Suc.) and Cycle Count Deviation (Cyc.). Since morse tapping has a fixed cycle count, we do not report its Cyc.
@@ 5.1 Experimental Setup
Real-world Setup. We conduct real-world experiments across diverse and heterogeneous robotic embodiments, including single-arm and dual-arm grippers (AgileX Piper), dexterous hands (BrainCO Revo2), and a humanoid platform (Untree G1). The visual observations are captured using an Intel RealSense L515 depth camera. There are six real-world cyclic manipulation tasks: block hammering, bottle shaking, drum beating, tire pumping, knife cutting, table cleaning. For each task, 50–150 teleoperated demonstrations are collected using Gello [wu2024gello] for gripper arms, TypeTele [lin2025typetelereleasingdexterityteleoperation] for dexterous hands, and OpenWBC [liu2025openwbc, ben2025homie] for the humanoid platform as shown in Figure 5. Additional details are provided in the supplementary material.
Simulation Setup We further evaluate our approach in simulation using the CycleManip benchmark and the RoboTwin 2.0 benchmark. In CycleManip simulation benchmark, two ARX-X5 robotic arms and a RealSense D435 head camera are employed, and eight cyclic tasks are performed as described in Section 4. For RoboTwin 2.0 benchmark, we follow the same experimental settings as in [RoboTwin2025Leaderboard, chen2025robotwin2] and evaluate four general manipulation tasks. More details are provided in the supplementary material.
@@ 5.2 Implementation Details
Training and Inference Details For our framework, the number of sampled high-overhead observations $K_{high}=6$ and the action horizon is 8. We employ DDIM [song2020ddim] as diffusion sampler with 100 training steps and 10 testing steps. For training, the number of epochs is set to 300 and the batch size is 128. The loss weights $\alpha=1$ and $\beta=0.1$. The initial learning rate is $2.0\times 10^{-4}$ with a cosine learning rate scheduler [loshchilov2016sgdr]. The experiments of our framework are conducted using PyTorch on a single RTX 4090 GPU. All real world experiments are inferred in a single RTX 4070 GPU. More details are provided in the supplementary material.
Baseline Reproduction We reproduce all baselines using the default settings of RoboTwin across all experiments. For the DP [chi2025dp] and DP3 [ze20243dp3] baselines, since they do not include a language encoder, we implement an identical language encoder within our framework for fair comparison. The Pi-0 [black2024pi0] and RDT [liu2024rdt] models are trained on a single H100, while DP and DP3 models are trained on a single RTX 4090. More details are provided in the supplementary material.
@@
task setting DP3 [ze20243dp3] w/o Task Ours
@@ 5.3 Comparison with State-of-the-Art Methods
place cans
plasticbox
Handover
Block
Pick Diverse
Bottles
Stamp
Seal
Place Bread
Basket
Open
Microwave
Turn
Switch
RDT [liu2024rdt] 6 45 2 1 10 37 35
Pi-0 [black2024pi0] 34 45 27 3 17 80 27
DP [chi2025dp] 40 10 6 2 14 5 36
DP3 [ze20243dp3] 48 70 52 18 26 61 46
@@
Table 3: The comparison results in general manipulation from RoboTwin 2.0 Benchmark [RoboTwin2025Leaderboard]. Our method not only yields benefits for cyclic tasks but also achieves good performance for general manipulation tasks.
bottle
shaking
rooller
rolling
carrot
cutting
dual-knife
chopping
Pi-0[black2024pi0] 19 14 8 1
@@ 5.5 Results in general manipulation
Table 3 presents a comprehensive comparison of various methods in general manipulation tasks from RoboTwin 2.0 benchmark [chen2025robotwin2]. Our method achieves the best performance across all tasks, far surpassing baselines such as RDT, Pi0, DP, and DP3. This consistent superiority demonstrates our approach’s remarkable robustness, generalization capability, and effectiveness in handling diverse robotic manipulation scenarios, validating its potential for advancing the state-of-the-art in this field.
@@ 5.7 Plug and Play Experiment
Table 4 demonstrates the effectiveness of plugging our method into other imitation policies, such as Vision-Language-Action model, Pi-0 [black2024pi0]. In implementation, we only sample the current frame for visual observations, then encode all past joints via a transformer before feeding them into the action expert of Pi-0. The results show that our method significantly improves the performance, highlighting its robustness and adaptability.
@@ 5.8 Efficiency Analysis
Table 5 compares the efficiency of our method with that of the baseline [chi2025dp] on the carrot-cutting task in our benchmark using a single RTX 4090 GPU. $\text{Time}_{\text{train}}$ denotes the per-step training time (1 diffusion step), and $\text{Time}_{\text{test}}$ denotes the inference time (10 diffusion steps). $\text{GPU}_{\text{train}}$ and $\text{GPU}_{\text{test}}$ indicate GPU memory usage. Our method achieves a higher success rate ($Suc.$) while incurring only marginally increased training and inference time, as well as slightly higher GPU memory consumption.

### v2

@@ 1 Introduction
The ability of a robot to autonomously take care of various tasks in our daily life is a long-term goal of the computer vision and robotics community [3, 46, 14, 20]. One key observation is that many tasks in the household involve repetitive and cyclic actions, such as dispensing several pumps of syrup or shaking a bottle until the contents are mixed [1, 11, 27]. These tasks require the robot to perform cyclic actions and stop at an expected time, which presents significant challenges: the robot needs to execute repetitive actions and complete them within the desired time frame according to the user command or task execution status. Previous work on robot manipulation has focused on general tasks [7, 23]; however, relatively little work has explored how to enable robots to perform cyclic tasks. Moreover, there is a lack of benchmarks that provide sufficient data and evaluation tools to assess the applicability of existing methods to cyclic tasks.
In this work, we explore this crucial task, CycleManip, the problem of robotic cyclic manipulation where a robot must accurately execute cyclic motions and stop at the correct moment. Such tasks are common in daily scenarios, but their inherently non-Markovian nature makes them challenging: the correct decision at any moment depends not only on the current observation, but also on the accumulated progress within the cycle. As a result, autonomous policies must model historical temporal dependencies and reason about the progression across multiple cycles[38].
However, traditional imitation policies typically rely on short observation windows for action prediction [3, 43, 4, 17, 32, 36, 33], which leads to failure in cyclic manipulation tasks. This is because short observations across cycles often appear similar, causing the model to confuse its decisions as shown in Figure 2 (a). For example, in a task that requires the robot to shake a bottle five times, the visual observation after each shaking step remains nearly identical, making it difficult for such models to infer whether the robot should continue or stop, since they cannot track how many shaking cycles have already been completed. An intuitive remedy is to expand the observation horizon, but doing so is costly: encoding and fusing high-dimensional visual observations at every timestep increases computation and latency.
@@
To support our framework, we present a cycle-based manipulation benchmark that provides a diverse set of simulated tasks with automated data generation and model evaluation tools. The benchmark is built on the RoboTwin 2.0 simulation platform [6], where we incorporate configurable cyclic action parameters into the data collection pipeline, enabling the generation of demonstrations with arbitrary numbers of repetitions and corresponding instructions. In addition, we develop an automated evaluation system in which an attempt is considered successful only if the policy not only completes the manipulation task but also performs the correct number of cycles.
@@ 2.1 Robotic Manipulation
Robotic manipulation is a fundamental challenge in robotics and is considered an essential cornerstone for achieving Artificial General Intelligence (AGI). Current dominant methodologies, such as Imitation Learning (IL) [7, 43, 46] and Vision-Language-Action (VLA) models [47, 17, 3, 13], excel at modeling complex data distributions. They effectively predict subsequent actions based on current observations, demonstrating strong performance in sequential tasks. However, these approaches falter in cyclic tasks. When faced with similar inputs, the model struggles to distinguish the current phase of the cycle, which can lead to it becoming trapped in infinite loops or terminating prematurely. To solve this problem, we propose the CycleManip framework to enable cyclic task manipulation in an end-to-end imitation manner by the effective perception and understanding of historical information.
@@ 2.2 Historical Modeling
Historical modeling is important for Large-Language-Models [37], Vision-Language-Navigation [45], video generation [40] and robot manipulation [29]. The historical information is used to address the limitations of short-term observations, by retrieval [42, 40], large kernels [8], memory caching [44, 21]. Recent robot manipulation methods integrate historical visual information for visual memory [21, 29, 31], with a primary focus on long-horizon tasks. In contrast, this paper addresses the more challenging domain of cyclic tasks and introduces a cost-aware sampling strategy to integrate historical low-overhead proprioceptive information.
@@ 2.3 Cyclic Task
Cyclic tasks are important in daily life and represent a critical challenge that needs to be addressed for robot deployment. The execution of cyclic tasks can significantly enhance efficiency [10, 9] and be crucial in environments such as factories, laboratories [1, 11, 27] and human-robot collaboration scenarios [16, 15]. Some traditional works implement cyclic tasks by control strategies, which may be limited by poor adaptability to dynamic environments. Recently, some deep learning-based methods have been attempted for simple cyclic tasks [41]; however, these methods are either limited to simple tasks with fixed cyclic times [34], limited to specific task scenarios [18, 12], or rely on external auxiliary models [5]. In this paper, we propose a method that enables diverse cyclic manipulation tasks with cyclic time control capability, in an end-to-end manner without leveraging any additional auxiliary models.
@@ 3.2 Effective Historical Perception
For high-overhead observations (e.g., point clouds for 3D-based methods and RGB images for 2D-based methods), we adopt a heuristic sampling strategy that selects past frames with a longer observation horizon while keeping the overall number of frames $K_{\text{high}}$ consistent with the original version. Specifically, given the first frame as 0 and the current frame as $t$, we perform right-sided binary sampling to collect $0.5\cdot K_{\text{high}}$ frames. And we apply exponential sampling from the latest frame $t$, collecting another $0.5\cdot K_{\text{high}}$ frames according to the rule $t-2^{k}$, where $k$ denotes the sampling index.
@@ 3.3 Effective Historical Understanding
Challenge of Historical Understanding. While expanding the observation horizon provides the model with broader perception, an overload of information introduces challenges in feature encoding and understanding. Relying only on imitation learning supervision typically fails to enable the model to truly understand the inherent temporal process of cyclic tasks, as shown in Figure 2 (b). For example, in a hammering task, the historical observations before each strike differ due to different motion histories, yet the supervision signal remains the same (conduct a hammering). This discrepancy may force the model to learn features that converage to a local optimum, making it difficult for the model to capture the temporal information necessary for cyclic behavior.
@@ 3.4 Framework Architecture
Given the user instruction lanlan and historical observation $\{o_{i}\}_{i=1}^{t}=\{o_{high},o_{low}\}_{i=1}^{t}$, we first employ a cost-aware sampling strategy to obtain sampled observation $\mathcal{H}_{h}\left(\{{o^{h}_{i}}\}_{i=1}^{t}\right)$ and $\mathcal{H}_{l}\left(\{{o^{l}_{i}}\}_{i=1}^{t}\right)$. The language feature $f_{lan}$ is encoded by CLIP encoder [26], the high observation feature $f_{h}$ is encoded by a point encoder [43]. For low-overhead observations, we employ a Transformer encoder [35] to obtain $f_{l}$, where the global features are derived from the CLS token, and local features are extracted using an MLP from the recent frames. Then we employ MLPs to fuse $f_{l}$ and $f_{h}$ obtain $f_{lh}$, which will be used for diffusion decision and auxiliary task prediction. Finally, we concatenate the language feature $f_{lh}$ and observation feature $f_{lh}$ as the condition feature for diffusion model. We employ film conditioning [25] to output the action prediction. The final loss function of the model is:
@@ 4 CycleManip Benchmark
To support our framework, we build a benchmark for cyclic manipulation tasks based on the RoboTwin 2.0 platform [6]. We collect 8 cyclic manipulation task environments for convenient data collection and policy evaluation, as shown in Figure 4. More details can be found in the supplementary material.
@@
block hammering bottle shaking roller rolling carrot cutting dual-knife chopping egg beating chemical mixing morse tapping
@@
Table 1: Performance comparison on various cyclic manipulation tasks. Our method outperforms all baselines in Success Rate (Suc. %) and Cycle Count Deviation (Cyc.). Since morse tapping has a fixed cycle count, we do not report its Cyc.
@@ 5.1 Experimental Setup
Real-world Setup. We conduct real-world experiments across diverse and heterogeneous robotic embodiments, including single-arm and dual-arm grippers (AgileX Piper), dexterous hands (BrainCO Revo2), and a humanoid platform (Untree G1). The visual observations are captured using an Intel RealSense L515 depth camera. There are six real-world cyclic manipulation tasks: block hammering, bottle shaking, drum beating, tire pumping, knife cutting, table cleaning. For each task, 50–150 teleoperated demonstrations are collected using Gello [39] for gripper arms, TypeTele [19] for dexterous hands, and OpenWBC [22, 2] for the humanoid platform as shown in Figure 5. Additional details are provided in the supplementary material.
Simulation Setup We further evaluate our approach in simulation using the CycleManip benchmark and the RoboTwin 2.0 benchmark. In CycleManip simulation benchmark, two ARX-X5 robotic arms and a RealSense D435 head camera are employed, and eight cyclic tasks are performed as described in Section 4. For RoboTwin 2.0 benchmark, we follow the same experimental settings as in [28, 6] and evaluate four general manipulation tasks. More details are provided in the supplementary material.
@@ 5.2 Implementation Details
Training and Inference Details For our framework, the number of sampled high-overhead observations $K_{high}=6$ and the action horizon is 8. We employ DDIM [30] as diffusion sampler with 100 training steps and 10 testing steps. For training, the number of epochs is set to 300 and the batch size is 128. The loss weights $\alpha=1$ and $\beta=0.1$. The initial learning rate is $2.0\times 10^{-4}$ with a cosine learning rate scheduler [24]. The experiments of our framework are conducted using PyTorch on a single RTX 4090 GPU. All real world experiments are inferred in a single RTX 4070 GPU. More details are provided in the supplementary material.
Baseline Reproduction We reproduce all baselines using the default settings of RoboTwin across all experiments. For the DP [7] and DP3 [43] baselines, since they do not include a language encoder, we implement an identical language encoder within our framework for fair comparison. The Pi-0 [3] and RDT [23] models are trained on a single H100, while DP and DP3 models are trained on a single RTX 4090. More details are provided in the supplementary material.
@@
task setting DP3 [43] w/o Task Ours
@@ 5.3 Comparison with State-of-the-Art Methods
place cans plasticbox Handover Block Pick Diverse Bottles Stamp Seal Place Bread Basket Open Microwave Turn Switch
RDT [23] 6 45 2 1 10 37 35
Pi-0 [3] 34 45 27 3 17 80 27
DP [7] 40 10 6 2 14 5 36
DP3 [43] 48 70 52 18 26 61 46
@@
Table 3: The comparison results in general manipulation from RoboTwin 2.0 Benchmark [28]. Our method not only yields benefits for cyclic tasks but also achieves good performance for general manipulation tasks.
bottle shaking rooller rolling carrot cutting dual-knife chopping
Pi-0[3] 19 14 8 1
@@ 5.5 Results in general manipulation
Table 3 presents a comprehensive comparison of various methods in general manipulation tasks from RoboTwin 2.0 benchmark [6]. Our method achieves the best performance across all tasks, far surpassing baselines such as RDT, Pi0, DP, and DP3. This consistent superiority demonstrates our approach’s remarkable robustness, generalization capability, and effectiveness in handling diverse robotic manipulation scenarios, validating its potential for advancing the state-of-the-art in this field.
@@ 5.7 Plug and Play Experiment
Table 4 demonstrates the effectiveness of plugging our method into other imitation policies, such as Vision-Language-Action model, Pi-0 [3]. In implementation, we only sample the current frame for visual observations, then encode all past joints via a transformer before feeding them into the action expert of Pi-0. The results show that our method significantly improves the performance, highlighting its robustness and adaptability.
@@ 5.8 Efficiency Analysis
Table 5 compares the efficiency of our method with that of the baseline [7] on the carrot-cutting task in our benchmark using a single RTX 4090 GPU. $\text{Time}_{\text{train}}$ denotes the per-step training time (1 diffusion step), and $\text{Time}_{\text{test}}$ denotes the inference time (10 diffusion steps). $\text{GPU}_{\text{train}}$ and $\text{GPU}_{\text{test}}$ indicate GPU memory usage. Our method achieves a higher success rate ($Suc.$) while incurring only marginally increased training and inference time, as well as slightly higher GPU memory consumption.
@@ 6 Conclusion
7 Acknowledgment
This work was supported partially by NSFC (92470202), Guangdong NSF Project (No. 2023B1515040025), Guangdong Key Research and Development Program (No. 2024B0101040004, No. 2025B0909020002). We would like to thank Jiangran Lyu for insightful and helpful discussions, and Pengxiang Liu for his work on robot hardware.